OpenStudy (anonymous):

Can someone help me with 4 problems I will give medals and become fans. I know I should re-post but it takes to long and gets annoying so I will not please do not ask. (Questions in comments) Please wait 2 minutes for the questions. (Thank you for helping me @iGreen and @radar on my other questions) Thank you for if you help me here before :)

OpenStudy (anonymous):

@iGreen should I multiply the two fractions

OpenStudy (igreen):

Area of a rectangle is \(l \times w\). \(6 \dfrac{2}{5} \times 4 \dfrac{1}{6}\) Convert into improper fractions: \(\dfrac{32}{5} \times \dfrac{25}{6}\rightarrow \dfrac{32 \times 25}{5 \times 6}\)

OpenStudy (anonymous):

I thought A am I corret

OpenStudy (igreen):

Yep, that's right.

OpenStudy (anonymous):

Can u tell me how to solve then I answer and its wrong you tell me the right one @iGreen @iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Change it into an improper fraction first. \(6 \dfrac{3}{4} \rightarrow \dfrac{27}{4}\) The reciprocal of a fraction is just switching the top and the bottom around. @starrhazel123

OpenStudy (anonymous):

So D @iGreen

OpenStudy (igreen):

No, switch the top and the bottom of 27/4 around, what do you get?

OpenStudy (anonymous):

NO WAIT A RIGHT :)

OpenStudy (anonymous):

A RIGHT

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

wait let me post next one @iGreen

OpenStudy (anonymous):

How do I solve this one

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen